OpenStudy (anonymous):

integral of sin^3(x)*cos^2(x) dx ?

OpenStudy (anonymous):

try \(u=\cos(x)\)

OpenStudy (anonymous):

u sub? or integration by parts?

OpenStudy (anonymous):

and rewrite \[\sin^3(x)\] as \[\sin(x)(1-\cos^2(x))\]

OpenStudy (anonymous):

u sub

OpenStudy (anonymous):

you get \[-\int(1-u^2)u^2du\] pretty much right away

OpenStudy (anonymous):

ok thanks! i think i got it from here

OpenStudy (anonymous):

yw